OpenStudy (anonymous):

DeAndre bought 15 party hats priced at 3 for $0.65 and 56 noisemakers priced at 7 for $1.25. What was the total cost of the hats and noisemakers, not including tax?

OpenStudy (nubeer):

one bundle of hats = 0.65 how many bundles = 15/3 = 5 so cost for hats = 5*0.65 = ? same way u have to calculate for noise makers.. can u do it?

OpenStudy (nubeer):

why? whats the hard part for u?

OpenStudy (anonymous):

all of it thats why i posted it.

OpenStudy (nubeer):

well then lets do it from start.. a bundle contains 3 hats and cost 0.65$ right?

OpenStudy (anonymous):

mmmhmmm

OpenStudy (nubeer):

so there are 15 hats.. and 1 bundle have 3 hats so how many bundles u have in total?

OpenStudy (anonymous):

5

OpenStudy (nubeer):

good.. so u have 5 bundles.. and 1 bundle cost 0.65 so 5 bundle will cost?

OpenStudy (anonymous):

3.25

OpenStudy (nubeer):

good.. now this is cost for ur hats.. now lets see for noise makers.. it says there is a bundle of 7 and cost 1.25.. first see how many bundle u will have if u have to buy 56 noise makers..

OpenStudy (anonymous):

i have no idea..

OpenStudy (nubeer):

u have to do same way u did above.. 1 bundle = 7noise makers. x bundles = 56 noise makers. (56/7)

OpenStudy (anonymous):

7x56?

OpenStudy (nubeer):

no.... 56/7 as this will give u number of bundles.. if u do 7x56 = 392 do u think 56 noise makers will make 392 bundles?

OpenStudy (anonymous):

ummmm yes

OpenStudy (nubeer):

lol do u think 392 bundles possible to just have 56 noise amkers just imagine... u have 56 pins.. and u have alot of boxes and in each box u can put just 7 pins( no less no more) so how many boxes u can fill?

OpenStudy (anonymous):

idk....

OpenStudy (nubeer):

just count atleast pick a box and put 7 pins in one box.. u can do like 56/7 this will tell u how many boxes will fill

OpenStudy (anonymous):

8

OpenStudy (nubeer):

ok good if 1 box is of 1.25$ how much 8 boxes will cost= ?

OpenStudy (anonymous):

9.80

OpenStudy (nubeer):

no .. 8*1.25 = 10 so noisemaker = 10$ hats = 3.25 cost = 3.25 + 10 = ? this would be your answer